# Voyager 1 Data Reveals Signatures of the Local Gas and Cosmic-Ray Source Distributions

### v1

@@ 3.4 Results
For all model configurations the light secondaries Li–Be–B are unaffected by the localised features for the CR or gas distributions. At high energies this is expected. But at the lower energies there is no significant effect either. This is showing that the production of these species, which are pure secondary for the 4022a scenario abundances that the 2D/baseline employs, is over a larger volume than the fairly nearby localised features (concentrations or deficits) in the gas or CR density or spectral fluctuations can affect the primary spectra.
@@
We note that the attenuation of the low-energy Carbon spectrum is still not sufficient to raise the B/C ratio to obtain complete consistency with the V1 data. This supports the contention of Cummings_2025 for a primary component contribution to the Boron spectrum.
@@
Figure 4 shows the elemental spectra for Z=11–19 (left) and Z=20–28 (right). The spectra $\lesssim$100 MeV nucleon-1 energies show sensitivity to the nearby ISM features of the models depending on their relative contributions via primary and secondary production channels. For example, Iron is mostly primary and shows a very strong effect for the low energy spectrum depending on source density hole size and gas model. (This is similar to the case for C and O in Fig. 2 that are also both mostly primary.) Meanwhile, the low-energy spectra of the mixed-origin species (primary+secondary) reflect the relative contributions from the primary abundance as well as respective fragmentation cross sections (Cummings_2025, Appendix B.4). Some of these elemental spectra, e.g., Vanadium have only a very small primary abundance as determined from the 2D model fit, and it is only for the S25/0 configuration that the spectrum appears significantly different from the baseline case.
@@
The calculated elemental ratios for He, O, and Fe for different setups are compared with V1 data in Figure 5. The V1 ratios are derived for each elemental combination by taking the intensity for each energy bin for whichever species has the least energy range coverage, then power-law interpolating at the respective mid-bin energies [^3] Following Cummings_2025 for given energy bin boundaries $E_{\rm max}$ and $E_{\rm min}$ this is ($E_{\rm min}+E_{\rm max}$)/2. to obtain the intensities of the other species. The uncertainties are likewise obtained using the same interpolation scheme combining the 1-$\sigma$ statistical uncertainties for each species (following the prescription from the note for Table C1 of Cummings_2025).
@@
For the left panel, which is showing the combinations with various hole sizes and the two gas models, the 2D fit (black solid) and S25/200 (cyan long dash dotted) are best matching to the data. The short and long black-dashed lines are the 3D equivalent with 100 pc and 200 pc radius holes in the source density distribution. The dotted, short dash dotted, and long dash dotted are calculations made with the S25 gas model and no hole/100 pc/200 pc radius hole in the source distribution, respectively.
@@
The shapes of the curves are defined by the interplay of several processes, such as the ionisation energy losses ($\propto Z^{2}$), fragmentation (roughly $\propto A^{0.6}$, but is quite different at low energies), and injection. The presence of the 100/200 pc-radius hole in the source distribution leads to the rise of the ratios toward low energies. This is mostly connected with the larger ionisation energy losses of the denominator species in the ratios, while the replenishment from the sources is absent. While the local gas number density in the vicinity of the Solar system in the S25 model is lower than in the 2D/3D equivalent model, the diffusion coefficient is also lower by $\sim$20%, which is a reflection of the effective large-scale gas distribution. Both effects reduce the ionisation energy losses and fragmentation of the denominator species, thus the blue curves are always lower than the black curves except at very low energies <<2–3 MeV nucleon-1 for setups with the holes in the source distribution (3D/100/200, S25/100/200). Note that the total fragmentation cross section for He is much smaller than that for the heavier species, O and Fe, see 2022ApJS..262...30P. For the latter, they are comparable even though the fragmentation cross section for Fe is larger than for O.
@@
Very fast ionisation energy losses are dominating at low energies, which leads to a universal spectral shape for all primary species, and this explains the flattening in the black dashed curves seen at low energies. The exact energy depends on $Z^{2}$, thus for the He/O and He/Fe ratios it is observed at lower energies than for the O/Fe ratio.
@@
The wavy shape of the ratios in the S25/0 setup appears due to a combination of the effects due to the ionisation energy losses, fragmentation, and injection. The rise of the fraction for the low energies ($\lesssim$10 MeV nucleon-1) is again due to the increased ionisation energy losses and fragmentation of the denominator species. The right panel shows the effect of changing the source density hole size, where we have introduced an intermediate hole size of 150 pc (double-dot dashed) and extreme one of 400 pc (triple-dot dashed) for the S25 configuration to better illustrate the progression. It is readily seen that the flattening shifts to even lower energies due to the lower local gas number density.
4 Discussion
@@
Our results show that low-energy CRs in the VLISM are jointly shaped by two distinct but physically linked forms of local inhomogeneity: the structured gas distribution, including the LB, and the corresponding deficit of recent CR accelerators within it. Prior work has typically examined only one of these factors at a time. Studies focused on source inhomogeneity using discrete, stochastic, or locally absent sources within a homogeneous ISM (e.g., 2021PhRvL.127n1101P), have shown that the distance and age of nearby SN remnants can modulate the spectra of low-energy primaries. Conversely, other investigations isolating effects of a local underdensity in the gas while retaining a smooth, Galaxy-wide source population (e.g., 2002A&A...381..539D) have mainly looked at effects on secondary CRs without considering energy losses. The picture that we have now of the LB and its surroundings suggests that it was excavated by SN explosions, leading to an interior rarified ISM and a highly structured distribution of neutral and molecular gas at its boundary. The SN remnants have long-since dispersed with the consequence that the space within the LB is also deficient in terms of CR injection. Our study provides the first explicit assessment of how these effects can act together to shape the low-energy CR spectra in the ISM.
@@
When the two forms of structure are combined, a consistent and physically motivated picture emerges. The S25 gas distribution captures the LB’s underdense interior and its irregular boundary of enhanced gas density, while the introduction of a $\sim$150–200 pc source region deficit about the Solar system models the absence of recent SN acitivity inside the cavity. Together, these components naturally reproduce the observed low-energy softening: particles originate at or just beyond the LB boundary, traverse regions of both reduced and elevated density, and accumulate energy losses in a manner consistent with the measured Voyager spectra. In this picture, the attenuation of primaries is neither dominated solely by propagation distance (as in a homogeneous ISM) nor artificially suppressed by a uniformly low density. Rather, it arises from the interaction between both structured CR injection and losses.
@@
The behaviour of secondary-dominated species is a further clue. Because the Li–Be–B production samples larger effective volumes, their fluxes remain relatively insensitive to localised features in either the gas or the source distribution, and their ratios vary only modestly across configurations that we have considered. However, as noted before when discussing Fig. 4, Cummings_2025 find evidence for a low-energy primary Boron contribution that would complicate interpretation of the B/C ratio in this picture. For this scenario (called 4022c in their work), we used the S25 gas distribution and followed the tuning procedure (Sec. 3.3) to normalise the configuration, and predict the corresponding B/C for hole sizes 0 and 150 pc. The results are shown in left panel of Fig. 6 together with those for the corresponding 4022a (pure secondary B) configuration with same hole sizes, as used earlier in this work. The energy range for this figure is reduced compared to earlier, because the models with primary Boron were fit only using the V1 data. The difference between configurations that include a primary Boron contribution and those that do not is as large, if not larger, than those associated with the inhomogeneous source and ISM distributions investigated already. Using this secondary/primary ratio both as a diagnostic for the LISM conditions as well as global propagation parameters may not be as reliable as conventionally taken.
@@
In Fig. 6 (right panel) we show the Be/O ratio that we suggest may be a more useful constraint for propagation scenarios. Data and uncertainties were obtained using the same interpolation procedure as employed for Fig. 5. Beryllium is a pure secondary, and the Be/O ratio is fairly insensitive to the low-energy effects seen for B/C with all scenarios consistent with the V1 data. At higher energies, the configurations with/without primary Boron may be distinguishable using higher-energy data from ACE-CRIS [^4] https://izw1.caltech.edu/ACE/ASC/level2/lvl2DATA_CRIS.html and AMS-02 (2021PhR...894....1A). This issue requires accurate Solar modulation models, e.g., HelMod (2020ApJS..250...27B), and will be addressed in a follow-up paper.
5 Summary
@@
The best agreement with V1 data is when there are no CR sources within $\sim$150–200 pc using the S25 gas reconstruction. For this scenario, the low-energy CR fluxes are coming from sources just beyond the LB boundary. Interplay between reduced energy losses in the underdense interior, enhanced losses in the region outside, and the proximity of the sources produce the observed spectra. The secondary-dominated species remain largely unaffected by these localised features, consequently the primary spectra carry the strongest imprint of nearby ISM structure.
The conclusions by Cummings_2025 about significant primary contribution to the observed Boron spectrum at low energies are supported by our work. This raises doubts in the widely accepted paradigm about Boron being 100% secondary in CRs and the derivation of propagation parameters from the B/C or B/O ratio.
@@
This work is partially funded via NASA grants 80NSSC22K0477, 80NSSC22K0718, and 80NSSC23K0169.

### v2

@@ III.4 Results
For all model configurations the light secondaries Li–Be–B are unaffected by the localised features for the CR or gas distributions. At high energies this is expected. But at the lower energies there is no significant effect either. This is showing that the production of these species, which are pure secondaries for the 4022a scenario abundances that the 2D/baseline employs, is over a larger volume than the fairly nearby localised features (concentrations or deficits) in the gas or CR density or that spectral fluctuations can affect the primary spectra.
@@
We note that the attenuation of the low-energy Carbon spectrum is still not sufficient to raise the B/C ratio to obtain complete consistency with the V1 data. This supports the contention of Cummings et al. (2025) for a primary component contribution to the Boron spectrum.
@@
Figure 4 shows the elemental spectra for Z=11–19 (left panel) and Z=20–28 (right panel). The spectra for $\lesssim$100 MeV nucleon-1 energies show sensitivity to the nearby ISM features of the models depending on their relative contributions via primary and secondary production channels. For example, Iron is mostly primary and shows a very strong effect for the low-energy spectrum depending on source density hole size and gas model. (This is similar to the case for C and O in Fig. 2 that are also both mostly primary.) Meanwhile, the low-energy spectra of the mixed-origin species (primary+secondary) reflect the relative contributions from the primary abundance as well as respective fragmentation cross sections (Cummings et al. 2025, Appendix B.4). Some of these elemental spectra (e.g., Vanadium) have only a very small primary abundance as determined from the 2D model fit, and it is only for the S25/0 configuration that the spectrum appears significantly different from the baseline case.
@@
The calculated elemental ratios for He, O, and Fe for different setups are compared with V1 data in Figure 5. The V1 ratios are derived for each elemental combination by taking the intensity for each energy bin for whichever species has the least energy range coverage and then using power-law interpolation at the respective mid-bin energies [^3] Following Cummings et al. (2025) for given energy bin boundaries $E_{\rm max}$ and $E_{\rm min}$ this is ($E_{\rm min}+E_{\rm max}$)/2. to obtain the intensities of the other species. The uncertainties are likewise obtained using the same interpolation scheme combining the 1-$\sigma$ statistical uncertainties for each species (following the prescription from the note for Table C1 of Cummings et al. 2025).
@@
For the left panel, which is showing the combinations with various hole sizes and the two gas models, the 2D fit (black solid) and S25/200 (cyan long dash dotted) are best matching to the data. The short and long black-dashed lines are the 3D equivalent with 100 and 200 pc radius holes, respectively, in the source density distribution. The dotted, short dash dotted, and long dash dotted are calculations made with the S25 gas model and no hole, 100 pc radius hole, and 200 pc radius hole in the source distribution, respectively.
@@
The shapes of the curves are defined by the interplay of several processes, such as the ionisation energy losses ($\propto Z^{2}$), fragmentation (roughly $\propto A^{0.6}$, but is quite different at low energies), and injection. The presence of the 100/200 pc radius hole in the source distribution leads to the rise of the ratios toward low energies. This is mostly connected with the larger ionisation energy losses of the denominator species in the ratios, while the replenishment from the sources is absent. While the local gas number density in the vicinity of the Solar system in the S25 model is lower than in the 2D/3D equivalent model, the diffusion coefficient is also lower by $\sim$20%, which is a reflection of the effective large-scale gas distribution. Both effects reduce the ionisation energy losses and fragmentation of the denominator species, thus the blue curves are always lower than the black curves except at very low energies <<2–3 MeV nucleon-1 for setups with the holes in the source distribution (3D/100/200, S25/100/200). Note that the total fragmentation cross section for He is much smaller than that for the heavier species, O and Fe, see Porter et al. (2022, Appendices F and G, Figs. 10 and 11). For the latter, they are comparable even though the fragmentation cross section for Fe is larger than for O.
@@
Very fast ionisation energy losses are dominating at low energies, which leads to a universal spectral shape for all primary species, and this explains the flattening in the black dashed curves seen at low energies. The exact energy depends on $Z^{2}$; thus, for the He/O and He/Fe ratios it is observed at lower energies than for the O/Fe ratio.
@@
The wavy shape of the ratios in the S25/0 setup appears due to a combination of the effects due to the ionisation energy losses, fragmentation, and injection. The rise of the fraction for the low energies ($\lesssim$10 MeV nucleon-1) is again due to the increased ionisation energy losses and fragmentation of the denominator species. The right panel shows the effect of changing the source density hole size, where we have introduced an intermediate hole size of 150 pc (double-dot dashed) and extreme one of 400 pc (triple-dot dashed) for the S25 configuration to better illustrate the progression. It is readily seen that the flattening shifts to even lower energies owing to the lower local gas number density.
IV Discussion
@@
Our results show that low-energy CRs in the VLISM are jointly shaped by two distinct but physically linked forms of local inhomogeneity: the structured gas distribution, including the LB, and the corresponding deficit of recent CR accelerators within it. Prior work has typically examined only one of these factors at a time. Studies focused on source inhomogeneity using discrete, stochastic, or locally absent sources within a homogeneous ISM (e.g., Phan et al. 2021), have shown that the distance and age of nearby SN remnants can modulate the spectra of low-energy primaries. Conversely, other investigations isolating effects of a local underdensity in the gas while retaining a smooth, Galaxy-wide source population (e.g., Donato et al. 2002) have mainly looked at effects on secondary CRs without considering energy losses. The picture that we have now of the LB and its surroundings suggests that it was excavated by SN explosions, leading to an interior rarified ISM and a highly structured distribution of neutral and molecular gas at its boundary (Zucker et al. 2022). The SN remnants have long since dispersed with the consequence that the space within the LB is also deficient in terms of CR injection. Our study provides the first explicit assessment of how these effects can act together to shape the low-energy CR spectra in the ISM.
@@
When the two forms of structure are combined, a consistent and physically motivated picture emerges. The S25 gas distribution captures the LB’s underdense interior and its irregular boundary of enhanced gas density, while the introduction of a $\sim$150–200 pc source region deficit about the Solar system models the absence of a recent and sufficiently strong hadronic CR injector inside the cavity. Together, these components naturally reproduce the observed low-energy softening: particles originate at or just beyond the LB boundary, traverse regions of both reduced and elevated density, and accumulate energy losses in a manner consistent with the measured Voyager spectra. In this picture, the attenuation of primaries is neither dominated solely by propagation distance (as in a homogeneous ISM) nor artificially suppressed by a uniformly low density. Rather, it arises from the interaction between both structured CR injection and losses.
The statistical likelihood of such a source deficit deserves consideration. Adopting canonical values for an order-of-magnitude expectation, we take a Galactic CR source rate of $\sim$1 per century and an effective “active” time window of $\Delta t\sim 10^{5}$ yr over which a single source can significantly influence the local sub-GeV/n nuclei spectra (e.g., Wiedenbeck et al. 1999). This implies $\sim$$10^{3}$ potentially relevant sources Galaxy-wide over $\Delta t$. If we approximate the spatial distribution as uniform over a star-forming disc of radius $R_{\rm disc}\sim 15$ kpc and half-thickness $Z_{\rm disc}\sim 0.1$ kpc (volume $V_{\rm disc}=\pi R_{\rm disc}^{2}(2Z_{\rm disc})\approx 1.4\times 10^{2}~{\rm kpc}^{3}$), then the expected number inside a sphere of radius $R$ is $\lambda(R)\approx(10^{3})\,V_{\rm sph}(R)/V_{\rm disc}$, with $V_{\rm sph}=4\pi R^{3}/3$. For $R=150$ pc, $V_{\rm sph}\approx 1.4\times 10^{-2}~{\rm kpc}^{3}$, giving $\lambda\simeq 0.10$ and hence $P(0)=e^{-\lambda}\simeq 0.90$. For $R=200$ pc, $V_{\rm sph}\approx 3.4\times 10^{-2}~{\rm kpc}^{3}$, giving $\lambda\simeq 0.24$ and $P(0)\simeq 0.79$. Even for a higher rate of 3 per century, probabilities remain significant at 0.74 and 0.49, respectively. Thus, from Poisson statistics alone, a local deficit of effective CR sources is plausible.
This statistical picture is supported by the observations of the distribution of Wolf–Rayet and OB stars. These stars are candidates for CR acceleration throughout their lives via powerful winds, in addition to being the progenitors of SNe (e.g., Moskalenko 2023). Crucially, recent 3D maps show that the star-forming regions containing these active and future CR sources are situated on the surface of the LB, expanding outward from its interior (Zucker et al. 2022). The fact that these massive stars are located at the boundary of the LB provides physical backing for our conclusion of a source-deficient volume within it.
While the evidence points to a lack of recent and future source activity within the LB, this must be reconciled with the known population of nearby pulsars. The ATNF Pulsar Catalogue lists a couple dozen objects within $\sim$200 pc. However, our finding applies specifically to hadronic CR sources. While pulsars are confirmed electron-positron accelerators, their capacity to accelerate nuclei to significant energies is less certain, although mechanisms for their production and release from the confining pulsar wind nebula have been investigated (e.g., Bednarek and Protheroe 2002; Bednarek and Bartosik 2004b, a).
Furthermore, due to large proper motions imparted by natal kicks, the current location of a pulsar can be far removed from the site of its parent SN. Given the age of the LB, the SN remnants that created it and any associated pulsars have long since dispersed or travelled significant distances. The presence of older, displaced pulsars is therefore not inconsistent with a local deficit of recent, powerful CR acceleration events. Consequently, our result does not contradict the existence of these pulsars but implies that either (i) these specific nearby objects are not significant contributors to the hadronic CR flux, or (ii) their contribution is subdominant to that from more distant sources, at least for the low-energy particles measured by V1.
@@
Meanwhile, the behaviour of secondary-dominated species is a further clue. Because the Li–Be–B production samples larger effective volumes, their fluxes remain relatively insensitive to localised features in either the gas or the source distribution, and their ratios vary only modestly across configurations that we have considered. However, as noted before when discussing Fig. 4, Cummings et al. (2025) find evidence for a low-energy primary Boron contribution that would complicate interpretation of the B/C ratio in this picture. For this scenario (called 4022c in their work), we used the S25 gas distribution and followed the tuning procedure (Sec. III.3) to normalise the configuration, and predict the corresponding B/C for hole sizes 0 and 150 pc. The results are shown in the left panel of Fig. 6 together with those for the corresponding 4022a (pure secondary B) configuration with the same hole sizes, as used earlier in this work. The energy range for this figure is reduced compared to earlier, because the models with primary Boron were fit only using the V1 data. The difference between configurations that include a primary Boron contribution and those that do not is as large, if not larger than, those associated with the inhomogeneous source and ISM distributions investigated already. Using this secondary/primary ratio both as a diagnostic for the LISM conditions and global propagation parameters may not be as reliable as conventionally taken.
@@
Figure 6 (right panel) shows the Be/O ratio that we suggest may be a more useful constraint for propagation scenarios. Data and uncertainties were obtained using the same interpolation procedure as employed for Fig. 5. Beryllium is a pure secondary, and the Be/O ratio is fairly insensitive to the low-energy effects seen for B/C with all scenarios consistent with the V1 data. At higher energies, the configurations with/without primary Boron may be distinguishable using higher-energy data from ACE-CRIS [^4] https://izw1.caltech.edu/ACE/ASC/level2/lvl2DATA_CRIS.html and AMS-02 (Aguilar et al. 2021). This issue requires accurate Solar modulation models, e.g., HelMod (Boschini et al. 2020), and will be addressed in a follow-up paper.
V Summary
@@
The best agreement with V1 data is when there is no dominant/recent source of CR nuclei within $\sim$150–200 pc (for the S25 gas reconstruction). For this scenario, the low-energy CR fluxes are coming from sources just beyond the LB boundary. While our modelling assumes a simplified source distribution outside this boundary, the fundamental conclusion that a combination of local gas structure and a local source deficit is necessary remains robust. Future work incorporating specific, powerful nearby catalogue sources like Vela and Geminga could further refine the exact parameters of this local void. Interplay between reduced energy losses in the underdense interior, enhanced losses in the region outside, and the proximity of the sources produce the observed spectra. The secondary-dominated species remain largely unaffected by these localised features; consequently, the primary spectra carry the strongest imprint of nearby ISM structure.
The conclusions by Cummings et al. (2025) about the significant primary contribution to the observed Boron spectrum at low energies are supported by our work. This raises doubts in the widely accepted paradigm about Boron being 100% secondary in CRs and the derivation of propagation parameters from the B/C or B/O ratio.
@@
This work is partially funded via NASA grants 80NM0018F0577, 80NSSC22K0477, 80NSSC22K0718, and 80NSSC23K0169.